Identify the center, vertices, and foci. Then sketch the graph.



((x + 3) ^ 2)/9 + (y ^ 2)/25 = 1

1
Expert's answer
2021-11-25T12:17:23-0500

Since the denominator of "y^2" is larger, the ellipse has a vertical major axis.

Standard form:

"\dfrac{y^2}{25}+\dfrac{(x+3)^2}{9}=1"


Center: "(-3, 0)"


"a^2=25, b^2=9,"

"c^2=a^2-b^2=25-9=16, c=4"

Foci: "(-3, -4), (-3, 4)"


Vertices: "(-3, -5), (-3, 5)"


Co-vertices: "(-6, 0), (0,0)"